# OPDR: Order-Preserving Dimension Reduction for Semantic Embedding of Multimodal Scientific Data

### v1

@@ Introduction
Incorporation into Scientific Data Management
We will first describe how the embedding vectors are generated from multimodal scientific data and then present how we implement the hypothetical function and incorporate it into popular dimension-reduction methods.
To convert scientific data into embedding vectors using CLIP (Contrastive Language-Image Pretraining), Vision Transformer (ViT), and BERT (Bidirectional Encoder Representations from Transformers), which are all transformer-based models, we begin by collecting and preparing the scientific data, which includes both textual data from HDF5 files and images in formats such as TIFF, PNG, or JPEG. The textual data is extracted by reading the Hierarchical Data Format (HDF5) files, where the dataset names and contents are concatenated into a single string representation. For instance, the HDF5 file used in our study contains a dataset with dimensions of (139, 139, 92) along with supporting metadata, providing a structured representation of the scientific data.
For the image data, each image file is read and converted to RGB format. The images, including results obtained at four different resolutions: HAADF, BF, ADF, and ABF, are found to have consistent dimensions of (369, 369) pixels. Additionally, a summary image with dimensions of (941, 905) pixels was included. Then, we generate embeddings for both textual and image data using the CLIP, ViT, and BERT models. For textual data, we utlized BERT model to generate embbedings, which processes the concatenated textual information. Image embeddings are produced by the ViT model, which processes the image files. Then, we utilized CLIP model to analyze both textual and image data.
The generated embeddings maintain their default dimensionality, which is specific to each model: BERT produces embeddings of 768 dimensions, while ViT also outputs 768-dimensional embeddings. For CLIP, the textual data is embedded via CLIP’s text encoder, resulting in 512-dimensional embeddings. Similarly, the image data is processed through CLIP’s image encoder, also yielding 512-dimensional embeddings. To create a comprehensive representation of the multimodal scientific data, we combined these embeddings by concatenating the 512-dimensional text and image vectors into a single, unified embedding vector of 1024 dimensions. All generated embeddings, derived from both textual data and image data at various resolutions, are collected and stored for further analysis.
Following the embedding vector extraction, we focus on the dimensionality reduction phase. We evaluated several mainstream techniques including PCA, and MDS, among which PCA consistently outperformed others in terms of maintaining the integrity of location information in our datasets. Our investigations revealed a notable correlation between the effectiveness of PCA in preserving the spatial relationships and the ratio $\frac{n}{m}$, where $n=dim(Y)$ and $m=\left|Y\right|$, of the target dimension to the number of samples used in the PCA process. We adopted various regression models to elucidate the relationship between the accuracy of preserved relative location information and the ratio $\frac{n}{m}$. These models facilitate the prediction of an optimal embedding vector dimension required to achieve a predetermined accuracy level, given a known number of samples, $m$.
@@
This section evaluates the effectiveness of the proposed method for KNN-preserved dimension reduction on multiple scientific data sets. To ensure the robustness of our method across various scenarios, we evaluate the method with three Transformer-based models for extracting embedding vectors, two popular dimension-reduction techniques, and three distance metrics. All results suggest that the proposed method is highly effective.
@@
The dataset used for evaluation were sourced from the Material Project  (The Materials Project Accessed 2024) and categorized into four distinct datasets: experimental Observable, stable, metal, magnetic. To be more specific, experimental observable dataset includes 33,990 data points, stable dataset consists of 48,884 data points, metal dataset has 72,252 data points, and magnetic dataset consists of 81,723 data points.
@@
For our experiments, we chose CloudLab (Duplyakin et al. 2019) as the testing platform. In this experiment, the specific machine specifications we selected are as follows: All of our experiments are carried out on a CloudLab machine with the following specifications: Node Type: c6420. The machine is equipped with two sixteen-core Intel Xeon Gold 6142 CPUs at 2.6 GHz, 384 GB ECC DDR4-2666 Memory, and two Seagate 1TB 7200 RPM 6G SATA hard drives. The network interface card (NIC) is a dual-port Intel X710 10GbE. We have installed the following libraries in conda environment: transformer, sklearn, torch, pandas, numpy, os, pickle, h5py, shutill, matlotlib, and pyprismatic.
@@
We first extracted embedding vectors from four datasets with distinct characteristics (Observable, Unstable, Metal, Non-magnetic) using the CLIP model. Subsequently, we computed the distances between vectors using the L2-norm and performed dimensionality reduction using PCA. To simplify notation, we let $n=dim(Y)$, denote the dimension of the lower-dimensional space. For these four groups of datasets, we further divided them into eight subsets each, with sample sizes $m$ where $m\in\{10,20,30,40,50,60,70,80\}$, to test whether the distribution patterns of data points were consistent across different sample sizes. Observations from figure 4 to figure 4 indicate a strong positive correlation between accuracy and the ratio of $n$ to $m$. As $n$ approaches $m$, accuracy initially increases rapidly and then slows down, converging to a stable value. Although the four different datasets exhibit highly similar data distributions, minor differences may be attributed to the inherent distribution characteristics and randomness of the samples.
@@
Through comparative analysis of the four data sets discussed above, we observe that our proposed OPDR algorithm performs consistently across multiple data sets and exhibits the anticipated data patterns, validating our initial hypotheses. In the remainder of this section, we will focus on the Observable Material dataset, exploring the effects of using different Transformer-based models, dimensionality reduction techniques, and various distance metrics.
@@
We continued our experiments on the ‘Observable’ dataset, merely changing the model used for extracting embedding vectors. We employed three distinct transformer-based models: BERT (Bidirectional Encoder Representations from Transformers), Vision Transformer (ViT), and CLIP (Contrastive Language-Image Pre-training) for processing and analyzing scientific data. Specifically, the BERT model was used to analyze text data extracted from HDF5 files, ViT was utilized for processing image data in TIFF and PNG formats, and CLIP was applied to analyze comprehensive multimodal datasets. This enabled us to achieve more holistic and unified representations of scientific data.
Through multiple experiments with these three models, we obtained data fit lines for each model to represent the relationship between accuracy and the ratio $n/m$. All three fit lines indicated that as $n$ approaches $m$, our accuracy first increases rapidly and then converges slowly. The results from the three different models were consistent with our initial hypotheses. Furthermore, we observed that the obtained fit lines were essentially overlapping, indicating that the choice of model does not alter the data structure. Utilizing these diverse transformer-based models, we were still able to ensure the accuracy and consistency of data relationships. Using different models does not change the data format; thus, in practical applications, it is essential to consider the properties of the data and its intended use to select the most appropriate model for experimentation.
Figure 5: Data Fitting for Different Embedding Models
Figure 5 shows that using different models does not alter much the accuracy. This implies that the closed-form function is applicable to a variety of neural network models. Therefore, in practical applications, one should consider the nature of the data and its intended use to select the most appropriate model for experimentation. The original data from which the regression line is calculated can be found in Figure 8 in the Appendix.
@@
The scientific data analyzed in this study exhibits extremely high dimensionality, often reaching up to hundreds of thousands of dimensions. Despite the application of advanced embedding techniques, the resulting vectors still contain several hundred to several thousand dimensions, posing considerable challenges for effective data analysis and interpretation. To tackle these issues, we employed two principal dimensionality reduction methods—Multidimensional Scaling (MDS) and Principal Component Analysis (PCA)—both renowned for their ability to efficiently reduce data complexity while maintaining the integrity of the relationships among data points. Given that the majority of contemporary dimensionality reduction techniques build on the foundational theories of PCA and MDS, our study specifically focused on these methods to evaluate their practical efficacy. By implementing MDS and PCA, we aimed to simplify the dataset in a way that preserves the relative positioning of data points, thereby facilitating more manageable and interpretable analysis.
By comparing the performance of PCA and MDS on the same dataset in Figure 6, we observe that PCA is more sensitive to changes in $n/m$ and can reach higher accuracy more quickly. Moreover, PCA can achieve a maximum accuracy of 100%, surpassing the peak accuracy attainable by MDS. This demonstrates PCA’s superior performance in preserving the relative positions of data points. Additionally, employing different dimensionality reduction techniques does not alter the overall data pattern; the relationship between accuracy and $n/m$ still aligns with our hypothesis.
Figure 6: Data Fitting for Different Dimension-Reduction Methods
From Figure 6, although the overall data pattern remains constant and both PCA and MDS align with our prior hypotheses, the choice of dimensionality reduction technique is found to have a noticeable effect on the fitting outcomes. In practical applications, it is crucial to weigh the specific contexts and experimental data at hand. This assessment will inform the selection of the most appropriate dimensionality reduction method to achieve the desired level of accuracy. The original data from which the regression line is calculated can be found in figure 9 in the Appendix.
Influence of Distance Metrics
Figure 7: Data Fitting for Different Distance Metrics
In this section, we explore the impact of using different distance metrics—namely L1 norm, L2 norm, and cosine distance—on the Observable dataset and CLIP model. The geometric meanings of these metrics are distinct:
  • L1 norm (Manhattan distance): Measures the sum of the absolute differences of their coordinates. It represents a very intuitive distance metric, similar to the shortest path between two points in a grid-like structure.
  • L2 norm (Euclidean distance): Calculates the straight-line distance between two points. It is the most commonly used distance metric, applicable in a wide range of natural and engineering problems.
  • Cosine distance: Measures the cosine of the angle between two vectors, reflecting the difference in direction rather than magnitude, often used to gauge similarity in text or high-dimensional data.
Experimental observations shown in Figure 7 revealed that as $n/m$ increases, the behaviors of L2 norm and cosine distance are similar, whereas L1 norm performs slightly worse than the other two. Despite the change in distance metrics not altering the overall data pattern—they still show the trend proposed in Eq (4)—the choice of distance metric remains crucial for specific scenarios.In practical use, experimenters should consider the application scenarios of their data to decide which distance calculation method to use. The original data from which the regression line is calculated can be found in figure 10 in the Appendix.
@@
This paper solves an open problem in multimodal scientific data management regarding the high dimensionality of embedding vectors generated by large language models. Theoretical results include new notions (both point-wise and space-level) to measure the closeness of two metric spaces in terms of their $k$-nearest neighbors and a closed-function to reveal the relationship between space dimensionality, space cardinality, and space similarity (or, accuracy of the dimension-reduction map). The closed-form function is implemented and incorporated into other subsystems, such as dimension-reduction methods (PCA, MDS), neural network models (CLIP, ViT, Bert), and distance metrics (Euclidean, cosine, Manhattan), which collectively enable a new dimension-reduction approach for efficiently manage and analyze multimodal scientific data.
Our future work will dive deeper into the theoretical foundation of the closed-form function among metric spaces’ dimensionality, cardinality, and accuracy of the dimension-reduction map; the fact that the proposed order-preserving measure is a well-defined measure may shed more insight than we originally thought. We also plan to investigate the extensibility of the proposed method for production vector database systems, such as PostgreSQL and pgvector.
Appendix A Additional Experimental Results
According to Figure 8, the data distributions obtained from the three Transformer-based models are highly similar. In contrast, the distribution from the BERT model is more dispersed, while the distributions from CLIP and ViT models are more concentrated. The results indicate that the CLIP model performs the most stably, with the overall patterns of the three models being consistent, aligning with our initial hypotheses. The fitting lines for the three groups of data points have been discussed previously in Figure 5.
Figure 8: Original Data for Neural Network Models
Figure 9 clearly shows the accuracy distributions for the ratio $n/m$ using two different dimensionality reduction techniques, PCA and MDS. Both PCA and MDS exhibit highly concentrated data points. Comparatively, PCA demonstrates higher accuracy as $n$ approaches $m$, indicating that PCA has stronger performance capabilities. Overall, the data patterns of both methods are similar, aligning with our initial hypotheses. The fitting lines for these two dimensionality reduction methods have already been discussed in figure 6.
Figure 9: Original Data for PCA and MDS
Figure 10 clearly illustrates the accuracy distributions for the ratio $n/m$ using different distance metrics. Data points generated by the L2-norm and cosine distance highly overlap, indicating similar performance of these metrics. In contrast, the L1-norm shows lower sensitivity and fails to reach the precision of other methods. Overall, the three datasets exhibit similar patterns: accuracy rapidly increases as $n$ approaches $m$ and then slowly converges to a peak. This observation aligns with our initial hypothesis. The fitting lines for the three sets of data points obtained using different distance metrics have been discussed in figure 7.
Figure 10: Original Data for Distance Metrics

### v2

@@ Introduction
Integration into Multimodal Data Retrieval
We will first describe how the embedding vectors are generated from multimodal data, specifically the image-text pair and then present how we implement the hypothetical function and incorporate it into popular dimension-reduction methods.
To leverage transformer-based models, including CLIP (Contrastive Language-Image Pretraining), Vision Transformer (ViT), and BERT (Bidirectional Encoder Representations from Transformers), for converting multimodal data into embedding vectors, we conducted experiments on three distinct datasets: materials science data, Flickr30k, and OmniCorpus. Each dataset consists of multimodal data, incorporating text from HDF5 files, natural language text, and images in TIFF, PNG, or JPEG formats. The experimental results demonstrate that our approach maintains strong applicability across different domains, including both scientific data and natural images.
The generated embeddings retain the default dimensionality of each respective model: BERT and ViT both produce 768-dimensional embeddings. For CLIP, the text and image encoders each output 512-dimensional vectors. We construct unified multimodal representations by directly concatenating embeddings from different modalities—for example, combining CLIP text and image vectors into a single 1024-dimensional vector.
In the case of audio-text data from the ESC-50 dataset, we generate embeddings using BERT (768D) for textual labels and PANNs CNN14 (Kong et al. 2020) (2048D) for audio signals. These are concatenated into 2816-dimensional joint vectors. All embeddings—spanning image, text, and audio modalities—are stored for subsequent dimensionality reduction and retrieval analysis.
Following the embedding vector extraction, we focus on the dimensionality reduction phase. We evaluated several mainstream techniques including PCA, and MDS, among which PCA consistently outperformed others in terms of maintaining the integrity of location information in our datasets. Our investigations revealed a notable correlation between the effectiveness of PCA in preserving the spatial relationships and the ratio
$\frac{n}{m}$, where $n=dim(Y)$ and $m=\left|Y\right|$, of the target dimension to the number of samples used in the PCA process. We adopted various regression models to elucidate the relationship between the accuracy of preserved relative location information and the ratio $\frac{n}{m}$. These models facilitate the prediction of an optimal embedding vector dimension required to achieve a predetermined accuracy level, given a known number of samples, $m$.
@@
This section evaluates the effectiveness of the proposed method for KNN-preserved dimension reduction on multiple data sets. To ensure the robustness of our method across various scenarios, we evaluate the method with several Transformer-based models for extracting embedding vectors, two popular dimension-reduction techniques, and three distance metrics. All results suggest that the proposed method is highly effective.
@@
The evaluation was conducted using seven datasets: four scientific datasets from the Materials Project (The Materials Project Accessed 2024), two natural image-text pair datasets (Flickr30k (Young et al. 2014) and OmniCorpus-037 CC (Li et al. 2024)), and one audio-text dataset (ESC-50 (Piczak 2015)).
Specifically, the observable, stable, metal, and magnetic datasets contain 33,990; 48,884; 72,252; and 81,723 data points, respectively. Flickr30k includes 31,014 image-text pairs, while OmniCorpus-037 CC provides 3,878,063 such pairs. The ESC-50 dataset comprises 2,000 environmental audio clips, each paired with a textual label describing the sound event (e.g., “dog barking”, “siren”, “rain”).
@@
Our experiments were carried out on the NSF-sponsored computing platform CloudLab (Duplyakin et al. 2019), specifically the d7525 nodes with the following specifications. The machine is equipped with two 16-core AMD EPYC 7302 CPUs at 3.00GHz, 128GB ECC Memory (8×16 GB 3200MT/s RDIMMs), and two 480 GB 6G SATA SSDs along with one 1.6 TB PCIe4 x4 NVMe SSD for storage. The network interface card (NIC) is a dual-port Mellanox ConnectX-6 DX 100Gb NIC, with one port supporting 200Gb connectivity for experimental use. Additionally, the machine is equipped with an NVIDIA 24GB Ampere A30 GPU.
We have installed the following libraries in conda environment: transformer, sklearn, torch, pandas, numpy, os, pickle, h5py, shutil, matplotlib, and pyprismatic, PIL.
@@
We first use the CLIP model to extract embedding vectors from six datasets, comprising four scientifically curated multimodal datasets (Observable, Unstable, Metal, Non-magnetic), two widely adopted vision-language benchmarks (Flickr30k and OmniCorpus), as well as the ESC-50 dataset (Piczak 2015) for audio-text representation. Subsequently, we computed the distances between vectors using the L2-norm and performed dimensionality reduction using PCA. To simplify notation, we let $n=dim(Y)$, denote the dimension of the lower-dimensional space. For these four groups of material datasets, we further divided them into eight subsets each, with sample sizes $m$ where $m\in\{10,20,30,40,50,60,70,80\}$. For the remaining two multimodal datasets, we divided them into five subsets each, with sample size $m$ where $m\in\{10,50,100,150,300\}$, to test whether the distribution patterns of data points remain consistent across different sample sizes. Observations from Figure 6 to Figure 6 indicate a strong positive correlation between accuracy and the ratio of $n$ to $m$. As $n$ approaches $m$, accuracy initially increases rapidly and then slows down, converging to a stable value. Although the four different datasets exhibit highly similar data distributions, minor differences may be attributed to the inherent distribution characteristics and randomness of the samples.
@@
Figure 5: Flicker30k
Figure 6: OmniCorpus
Through comparative analysis of the six data sets discussed above, we observe that our proposed OPDR algorithm performs consistently across multiple data sets and exhibits the anticipated data patterns, validating our initial hypotheses. In the remainder of this section, we will focus on the Observable Material dataset, exploring the effects of using different Transformer-based models, dimensionality reduction techniques, and various distance metrics.
@@
To evaluate the impact of different embedding models, we conducted further experiments on the Observable dataset, alongside Flickr30k and OmniCorpus from the multimodal datasets, using different models for extracting embedding vectors. We employed three distinct transformer-based models—BERT (Bidirectional Encoder Representations from Transformers), Vision Transformer (ViT), and CLIP (Contrastive Language-Image Pre-training)—to process and analyze the datasets. This enabled us to obtain more holistic and unified representations across different data domains.
Through extensive experiments with these three models, we derived data fit lines to represent the relationship between accuracy and the ratio $n/m$. All three fit lines consistently indicated that as $n$ approaches $m$, accuracy initially increases rapidly and then gradually converges. The results from the three different models aligned with our initial hypotheses. Additionally, for datasets with structured relationships—such as material-related datasets—the obtained fit lines were nearly overlapping, suggesting that the choice of model does not significantly alter the underlying data structure. In contrast, for the Flickr30k and OmniCorpus datasets, differences between models were more pronounced, though the overall trend still followed the expected logarithmic pattern. This indicates that while different transformer-based models may introduce variations in representation, they maintain the fundamental structure of the data. These findings reinforce that, regardless of the chosen model, the core data relationships remain stable. Thus, in practical applications, selecting the most appropriate model should be guided by the characteristics of the dataset and the specific task requirements, rather than concerns over fundamental data integrity.
Figure 7: Data Fitting for Embedding Models on Material Dataset
Figure 8: Data Fitting for Embedding Models on Flicker dataset
Figure 9: Data Fitting for Embedding Models on OmniCorpus dataset
Figure 10: Data Fitting for Dimension Reduction on Material Dataset
Figure 11: Data Fitting for Dimension Reduction on flicker dataset
Figure 12: Data Fitting for Dimension Reduction on OmniCorpus dataset
Figure 12 to Figure 12 show that using different models does not influence much the accuracy. This implies that the closed-form function is applicable to a variety of neural network models. Therefore, in practical applications, one should consider the nature of the data and its intended use to select the most appropriate model for experimentation.
@@
The datasets analyzed in this study exhibit extremely high dimensionality, often reaching hundreds of thousands of dimensions. Even after applying advanced embedding techniques, the resulting vectors typically retain several hundred to several thousand dimensions, presenting significant challenges for effective analysis and interpretation. To address these issues, we employed two principal dimensionality reduction methods—Multidimensional Scaling (MDS) and Principal Component Analysis (PCA)—both widely recognized for their ability to efficiently reduce data complexity while preserving the structural integrity of data relationships. As many modern dimensionality reduction techniques are fundamentally derived from PCA and MDS, our study specifically focused on these methods to assess their practical effectiveness.
By comparing the performance of PCA and MDS on the same dataset in Figure 12, we found that PCA is more sensitive to changes in $n/m$ and converges to higher accuracy more quickly. Additionally, in the materials science dataset, PCA achieves a peak neighborhood preservation accuracy of 100%, surpassing the highest accuracy attainable by MDS. In the other two multimodal datasets, PCA also demonstrates greater precision compared to MDS, further validating its superiority. Moreover, while different dimensionality reduction techniques may affect specific numerical values, they do not alter the overall data pattern; the relationship between accuracy and $n/m$ remains consistent with our hypothesis.
From Figure 12 to Figure 12, although the overall data pattern remains constant and both PCA and MDS align with our prior hypotheses, the choice of dimensionality reduction technique noticeably affects the fitting outcome. In practical applications, it is crucial to weigh the specific contexts and experimental data at hand. This assessment will inform the selection of the most appropriate dimensionality reduction method to achieve the desired level of accuracy.
@@
This paper proposes a new method, namely Order-Preserving Dimension Reduction (OPDR), to address the challenge of high-dimensional embeddings in multimodal data analytics, particularly those generated by large language models. From a theoretical perspective, this work introduces new concepts (both point-wise and space-level) to measure the closeness of two metric spaces in terms of their $k$-nearest neighbors and a closed-form function to characterize the relationship among dimensionality, cardinality, and similarity (i.e., the accuracy of the dimension-reduction map). The closed-form function is incorporated into a practical ecosystem by complementing other system components, such as dimension-reduction methods (PCA, MDS) and models (CLIP, ViT, Bert). Extensive evaluation demonstrates both the effectiveness and practical utility of OPDR.
Our future work will further explore the theoretical foundation of the closed-form function, particularly the relationship among metric spaces’ dimensionality, cardinality, and the accuracy of the dimension-reduction map. The fact that the proposed order-preserving measure is a well-defined metric may provide deeper insights. Additionally, we plan to investigate the extensibility of our method in production vector database systems, such as PostgreSQL and pgvector, to support large-scale retrieval tasks. Beyond multimodal retrieval, we aim to explore potential applications in medical imaging, where high-dimensional embeddings from MRI (Chamberland et al. 2019; Sridhar et al. 2022), CT scans, and vital records could benefit from dimension reduction while preserving similarity relationships—enabling efficient, accurate, and scalable content-based retrieval for medical diagnosis.